# AlphaDecay: Module-wise Weight Decay for Heavy-Tailed Balancing in LLMs

### v2

@@ 4.3 Analysis
To investigate the impact of various HT-SR metrics on regulating weight decay during model training, we conducted ablation studies comparing these metrics. While prior work has primarily utilized GradNorm [31; 20; 43] and FrobeniusNorm [12; 16] as indicators for adjusting weight decay, our study further evaluates additional metrics, including PL_Alpha_Hill and SpectralNorm, under the same experimental settings. Results in Figure 6 show that most HT-SR metrics outperform the uniform baseline, while PL_Alpha_Hill achieves the lowest perplexity (lower the better) among all evaluated methods.
@@
In our proposed framework, the HT-SR metric PL_Alpha_Hill is employed to guide the projection-based adjustment of weight decay during training. Since PL_Alpha_Hill is derived through PL fitting, and the choice of fitting method can influence both computational efficiency and the final training effectiveness, we conduct an ablation study to systematically assess its impact. Figure 7 presents a comparative analysis of three PL fitting methods— Goodness-of-fit [1; 30; 6], Fix-finger[44], and Median[47] —across multiple weight decay values. Across all settings, Median not only ensures optimal training performance but also notably decreases computation time compared to the other approaches, making it the preferred choice for PL fitting within our method.
LLaMa-60M LLaMa-135M
Weight Decay 0.1 0.05 0.01 0.1 0.05 0.01
WD=0 32.73 24.39
Uniform 31.95 32.31 32.66 23.32 23.81 24.28
AWD 32.58 32.67 32.67 24.30 24.35 24.41
Adadecay 31.88 32.25 32.58 23.18 23.62 24.21
AlphaDecay 31.20 31.65 32.45 22.66 23.04 23.98
@@
4.4 LLM Pre-training with AdamW
4.5 Dependent t-test for paired samples
[…]
AlphaDecay 22.77 $\pm$ 0.02 22.54 $\pm$ 0.03 23.44 $\pm$ 0.02
4.6 Evaluation with vision transformers
To evaluate the impact of different weight decay scheduling strategies on vision models, we train ViT-tiny on ImageNet-1K for 120 epochs using the Adam optimizer and ConvNeXt training configurations [24], applying each scheduling method as described in Table 5.
[…]
Weight Decay Unifrom AWD AdaDecay AlphaDecay
0 65.12
5e-6 66.97 65.21 65.86 67.40
1e-6 66.41 64.98 66.26 67.73
The results in Table 5 show that AlphaDecay consistently outperforms other strategies, with more pronounced accuracy improvements over Uniform and AWD. These findings indicate that the benefits of dynamic weight decay scheduling, particularly AlphaDecay, generalize well to vision transformer models.
@@ A.1 Architecture
To ensure reproducibility and consistency with prior research (Zhao et al. [2024], Lialin et al. [2023]), we adopt the LLaMa architectural specifications and pre-training hyperparameters as detailed in Table 6. All model variants are trained with a uniform maximum sequence length of 256, a batch size of 512, and an aggregate of 13K tokens per batch.
@@ A.2 Hyperparameter settings for reproducing our figures
Uniform LLaMa 60M 1e-5 33.002 - Uniform LLaMa 135M 1e-5 23.111 -
@@
Scaling Ratio $(s_{1},s_{2})$
1e-5 32.166 8.73$\pm$0.30 32.231 7.90$\pm$0.33 31.628 1.67$\pm$0.01 (0.67,3)
5e-6 32.436 10.62$\pm$0.04 32.381 9.22$\pm$0.47 31.614 1.69$\pm$0.01 (0.67,5)
LLaMa -60M 1e-6 32.993 8.28$\pm$0.03 33.059 7.80$\pm$0.67 32.703 1.66$\pm$0.01 (0.67,5)
1e-5 22.937 24.86$\pm$0.11 23.004 21.53$\pm$0.83 23.004 4.67$\pm$0.02 (0.67,3)
5e-6 22.937 24.43$\pm$0.11 23.090 22.00$\pm$0.86 22.588 4.68$\pm$0.03 (0.67,5)
LLaMa -135M 1e-6 23.924 24.64$\pm$0.13 24.058 21.90$\pm$0.80 23.448 4.63$\pm$0.01 (0.67,5)
@@ A.3 Assignment Function Formulas
Appendix B Varying assignment function hyperparameters

### v3

@@ 4.4 Analysis
To investigate the impact of various HT-SR metrics on regulating weight decay during model training, we conducted ablation studies comparing these metrics. While prior work has primarily utilized GradNorm [nakamura2019adaptive; kosson2023rotational; xie2023overlooked] and FrobeniusNorm [ghiasi2023improving; ishii2017layer] as indicators for adjusting weight decay, our study further evaluates additional metrics, including PL_Alpha_Hill and SpectralNorm, under the same experimental settings. Results in Figure 6 show that most HT-SR metrics outperform the uniform baseline, while PL_Alpha_Hill achieves the lowest perplexity (lower the better) among all evaluated methods.
@@
In our proposed framework, the HT-SR metric PL_Alpha_Hill is employed to guide the projection-based adjustment of weight decay during training. Since PL_Alpha_Hill is derived through PL fitting, and the choice of fitting method can influence both computational efficiency and the final training effectiveness, we conduct an ablation study to systematically assess its impact. Figure 7 presents a comparative analysis of three PL fitting methods— Goodness-of-fit [alstott2014powerlaw; martin2021predicting; clauset2009power], Fix-finger[yang2023test], and Median[zhou2023temperature] —across multiple weight decay values. Across all settings, Median not only ensures optimal training performance but also notably decreases computation time compared to the other approaches, making it the preferred choice for PL fitting within our method.
@@ A.1 Architecture
To ensure reproducibility and consistency with prior research (zhao2024galore, lialin2023relora), we adopt the LLaMa architectural specifications and pre-training hyperparameters as detailed in Table 6. All model variants are trained with a uniform maximum sequence length of 256, a batch size of 512, and an aggregate of 13K tokens per batch.
@@ A.2 Hyperparameter settings for reproducing our figures
Uniform
LLaMa
60M
1e-5 33.002 - Uniform
LLaMa
135M
1e-5 23.111 -
@@
Scaling Ratio
$(s_{1},s_{2})$
[…]
LLaMa
-60M
[…]
LLaMa
-135M
[…]
@@ A.3 Assignment Function Formulas
A.4 Derivation of Formula 2
We provide the derivation for estimating the power-law exponent $\alpha$ from empirical singular value data, which is central to our approach. We assume the empirical distribution follows a power-law (Pareto) distribution:
$p(x)=cx^{-\alpha}$
For normalization over $x\geq x_{min}$:
$\int_{x_{\min}}^{\infty}p(x)\,dx=1\implies c=(\alpha-1)x_{\min}^{\alpha-1}$
Thus, the probability density function (PDF) becomes:
$p(x)=(\alpha-1)x_{\min}^{\alpha-1}x^{-\alpha}=\frac{\alpha-1}{x_{\min}}\left(\frac{x}{x_{\min}}\right)^{-\alpha}$
Given a set of observed data $x_{1},x_{2},\dots,x_{n}$ with $x_{i}\geq x_{min}$, the likelihood function is:
$p(x_{1},x_{2},\dots,x_{n})=\prod_{i=1}^{n}p(x_{i})=\prod_{i=1}^{n}\frac{\alpha-1}{x_{\min}}\left(\frac{x_{i}}{x_{\min}}\right)^{-\alpha}$
The log-likelihood is therefore:
$\mathcal{L}=\sum_{i=1}^{n}\left[\ln(\alpha-1)-\ln x_{\min}-\alpha\ln\frac{x_{i}}{x_{\min}}\right]$
To obtain the maximum likelihood estimate, we set the derivative of $\mathcal{L}$ with respect to $\alpha$ to zero:
$\frac{\partial\mathcal{L}}{\partial\alpha}=0\implies\alpha=1+n\left[\sum_{i=1}^{n}\ln\frac{x_{i}}{x_{\min}}\right]^{-1}$
This result is known as the standard Hill estimator, and we denote the fitted exponent as PL_Alpha_Hill.
Appendix B More experiments
B.1 LLM Pre-training with AdamW
LLaMa-60M LLaMa-135M
Weight Decay 0.1 0.05 0.01 0.1 0.05 0.01
WD=0 32.73 24.39
Uniform 31.95 32.31 32.66 23.32 23.81 24.28
AWD 32.58 32.67 32.67 24.30 24.35 24.41
Adadecay 31.88 32.25 32.58 23.18 23.62 24.21
AlphaDecay 31.20 31.65 32.45 22.66 23.04 23.98
B.2 Dependent t-test for paired samples
Table 13: (Dependent t-test with Adam.) Each method (Uniform, AWD, AdaDecay, and AlphaDecay) is evaluated by conducting six repeated experiments with random seeds $\{$ 5, 6, 7, 8, 9, 10 }\}. Validation perplexity is reported as mean ± standard deviation. For each weight decay setting, a dependent t-test for paired samples is performed, comparing AlphaDecay against Uniform, AWD, and AdaDecay, respectively. The resulting p-values are presented alongside perplexity scores.
AlphaDecay 22.77 $\pm$ 0.02 22.54 $\pm$ 0.03 23.44 $\pm$ 0.02
B.3 Varying assignment function hyperparameters